# Finite-Time Analysis of Simultaneous Double Q-learning

### v1

@@ 3.4 Assumption and Definition
Throughout, we make the following standard assumptions, which are widely adopted in the reinforcement learning (RL) literature. We consider the scenario where the data is collected by the RL agent following a behavior policy $\beta$, which is the policy that the agent actually uses to interact with the environment and collect experiences. Moreover, we assume that the state-action pair at each time step is sampled i.i.d. from the stationary state distribution $p$ and the behavior policy $\beta$, respectively. Under these assumptions, the joint state-action distribution can be represented by
$d(s,a)=p(s)\beta(a|s),\quad(s,a)\in S\times A.$
@@
Assumption 3.1.
$d(s,a)>0$ holds for all $s\in S,a\in A$.
Assumption 3.2.
The step-size is a constant $\alpha\in(0,1).$
Assumption 3.3.
We have
$\max_{(s,a,s^{\prime})\in S\times A\times S^{\prime}}\lvert r(s,a,s^{\prime})\rvert=R_{\text{max}}\leq 1.$
Assumption 3.4.
The initial iterate $Q_{0}$ satisfies $\lVert Q_{0}\rVert_{\infty}\leq 1.$
3.1 is required for sufficient explorations, and 3.3 and 3.4 are introduced without loss of generality and for simplicity of the analysis. The following quantities will be frequently used in this paper; hence, the following notations are defined for convenience.
@@
$d_{\text{max}}\coloneq\max_{(s,a)\in S\times A}d(s,a)\in(0,1).$
@@
$d_{\text{min}}\coloneq\min_{(s,a)\in S\times A}d(s,a)\in(0,1).$
@@
$\rho\coloneq 1-\alpha d_{\text{min}}(1-\gamma).$ (4)
Under 3.2, the decay rate satisfies $\rho\in(0,1)$. Throughout the paper, we will use the following compact notations for dynamical system representations:
$P=\begin{bmatrix}P_{1}\\ \vdots\\ P_{\lvert A\rvert}\end{bmatrix},R=\begin{bmatrix}R_{1}\\ \vdots\\ R_{\lvert A\rvert}\end{bmatrix},Q=\begin{bmatrix}Q(\cdot,1)\\ \vdots\\ Q(\cdot,\lvert A\rvert)\end{bmatrix},$
$D_{a}=\begin{bmatrix}d(1,a)&&\\ &\dots&\\ &&d(\lvert S\rvert,a)\end{bmatrix},\quad D=\begin{bmatrix}D_{1}&&\\ &\dots&\\ &&D_{\lvert A\rvert}\\ \end{bmatrix},$
where $P_{a}=P(\cdot\lvert a,\cdot)\in\mathbb{R}^{\lvert S\rvert\times\lvert S\rvert}$, $Q(\cdot,a)\in\mathbb{R}^{\lvert S\rvert}$, $a\in\textit{A}$, and $R_{a}(s)\coloneqq\mathbb{E}\left[r(s,a,s^{\prime})\lvert s,a\right]$. Note that $P\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert\times\lvert S\rvert}$, $R\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert}$, $Q\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert}$, and $D\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert\times\lvert S\rvert\lvert A\rvert}$. Using these notations, $Q$-function is encoded as a single vector $Q\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert}$, which enumerates $Q(s,a)$ for all $s\in S$ and $a\in A$. The single value $Q(s,a)$ can be written as $Q(s,a)=(e_{a}\otimes e_{s})^{T}Q$, where $e_{s}\in\mathbb{R}^{\lvert S\rvert}$ and $e_{a}\in\mathbb{R}^{\lvert A\rvert}$ are $s$-th basis vectors (all components are 00 except for the $s$-th component which is 1) and $a$-th basis vectors, respectively, and $\otimes$ represents the Kronecker product. Under 3.2, $D$ is a non-singular diagonal matrix with strictly positive diagonal elements. For any stochastic policy, $\pi:S\rightarrow\Delta_{\lvert A\rvert}$, where $\Delta_{\lvert A\rvert}$ is the set of all probability distributions over A, let us define the matrix
$\Pi^{\pi}\coloneqq\begin{bmatrix}\pi(1)^{T}\otimes e_{1}^{T}\\ \pi(2)^{T}\otimes e_{2}^{T}\\ \vdots\\ \pi(\lvert S\rvert)^{T}\otimes e_{\lvert S\rvert}^{T}\\ \end{bmatrix}\in\mathbb{R}^{\lvert S\rvert\times\lvert S\rvert\lvert A\rvert,}$ (5)
where $e_{s}\in\mathbb{R}^{\lvert S\rvert}$. Then, it is well known that $P\Pi^{\pi}\in{\mathbb{R}}^{|{\cal S}||{\cal A}|\times|{\cal S}||{\cal A}|}$ is the transition probability matrix of the state-action pair under policy $\pi$. If we consider a deterministic policy, $\pi:{\cal S}\to{\cal A}$, the stochastic policy can be replaced with the corresponding one-hot encoding vector $\vec{\pi}(s):=e_{\pi(s)}\in\Delta_{|{\cal A}|},$ where $e_{a}\in{\mathbb{R}}^{|{\cal A}|}$, and the corresponding action transition matrix is identical to (5) with $\pi$ replaced with $\vec{\pi}$. For any given $Q\in{\mathbb{R}}^{|{\cal S}||{\cal A}|}$, denote the greedy policy w.r.t. $Q$ as $\pi_{Q}(s):=\operatorname*{arg\,max}_{a\in{\cal A}}Q(s,a)\in{\cal A}$. We will use the following shorthand frequently: $\Pi_{Q}:=\Pi^{\pi_{Q}}$.
The boundedness of $Q$-learning iterates [23] plays an important role in our analysis.
@@
$\lVert Q_{k}\rVert_{\infty}\leq Q_{\text{max}}=\frac{\text{max}{\{R_{\text{max}},\text{max}_{(s,a)\in S\times A}Q_{0}(s,a)\}}}{1-\gamma}.$
From Assumptions 3.3 and 3.4, we can easily see that $Q_{\max}\leq\frac{1}{1-\gamma}$.
@@ 4.1 Algorithm
$Q_{k+1}^{A}(s_{k},a_{k})$ $=Q_{k}^{A}(s_{k},a_{k})+\alpha_{k}\{r_{k}+\gamma Q_{k}^{A}(s_{k+1},\text{argmax}_{a\in B}Q_{k}^{B}(s_{k},a))-Q_{k}^{A}(s_{k},a_{k})\},$
$Q_{k+1}^{B}(s_{k},a_{k})$ $=Q_{k}^{B}(s_{k},a_{k})+\alpha_{k}\{r_{k}+\gamma Q_{k}^{B}(s_{k+1},\text{argmax}_{a\in A}Q_{k}^{A}(s_{k},a))-Q_{k}^{B}(s_{k},a_{k})\}.$ (6)
@@
The first difference between the original double $Q$-learning and SDQ is the role of each $Q$-estimator in the update. In the original double $Q$-learning, an optimal action is selected from the same $Q$-estimator, and it employs the other $Q$-estimator for bootstrapping. On the other hand, in the proposed version, an optimal action is selected from the other $Q$-estimator, and it employs the same $Q$-estimator for bootstrapping. This modification enables applying the switching system framework from [12], overcoming the difficulty caused by the switched order of $Q_{A}$ and $Q_{B}$ in original double $Q$-learning, while retaining its advantage of reducing overestimation bias.
@@
The other difference is in the Bernoulli variable. Unlike the standard double $Q$-learning, which uses a Bernoulli variable for the $Q$-estimator selection, the modified version updates the two $Q$-estimators synchronously, which can potentially speed up the convergence. However, we note that our analysis can also include the Bernoulli random selection as in the original form without major changes in the finite-time error analysis. Besides, a potential issue that arises by eliminating the random $Q$-estimator selection is that if initially $Q^{A}_{0}=Q^{B}_{0}$, then $Q^{A}_{k}=Q^{B}_{k}$ for all $k\geq 0$. In this case, (4.1) is reduced to the standard $Q$-learning. To bypass the issue for implementation, one simple approach is to randomly initialize $Q^{A}_{0}$ and $Q^{B}_{0}$ so that $Q^{A}_{0}\neq Q^{B}_{0}$.
@@
To demonstrate its effect, let us consider the example in Figure 1(left) (adopted from [1], Ch. 6.7). As observed in the result, SDQ initially suffers from overestimation of $Q$-values due to the random initialization of $Q$-estimators with uniform distribution. However, this bias is quickly mitigated. And it achieves similar result to the original double $Q$-learning as shown in Figure 1(right).
[…]
@@ 4.2 Experiment
In this section, we empirically compare SDQ with original double $Q$-learning on the environments, FrozenLake, CliffWalking, and Taxi in OpenAI Gym, and demonstrate that SDQ outperforms the standard double $Q$-learning. We utilize the epsilon-greedy exploration with the constant step-size $\alpha=0.01$ and discount factor $\gamma=0.99$ for both algorithms. $Q$-estimators for SDQ are initialized randomly with the uniform distribution, and intel core i7-11800h is used for all experiments. To account for the sparse rewards in FrozenLake, we evaluate the average episodic reward after applying a moving window of size $100$ for each episode. This approach smooths the reward signal, as the agent only receives a $+1$ reward upon successful completion. For all experiments, the results are averaged over $30$ independent runs. For comparative analysis, the baseline algorithms from https://github.com/ShangtongZhang/reinforcement-learning-an-introduction/tree/master/chapter06 are borrowed. As can be seen from Figure 2, SDQ exhibits better performances over the standard double $Q$-learning.
@@
Figure 2: Comparison of experiment results: Simultaneous Double $Q$-Learning (SDQ) vs. Double $Q$-Learning

### v2

@@ 3.4 Assumption and Definition
Throughout, we make the following standard assumptions, which are widely adopted in the RL literature. We consider the scenario where the data samples are generated from an RL agent interacting with the environment under a fixed behavior policy $\beta$. At each iteration, the state–action pair $(s,a)$ is assumed to be drawn independently from the stationary state distribution $p$ and the behavior policy $\beta$, which leads to the following joint distribution:
$d(s,a)=p(s)\beta(a|s),\quad(s,a)\in\mathcal{S}\times\mathcal{A}.$
@@
Assumption 1.
(Sufficient exploration) $d(s,a)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$.
Assumption 2.
(Constant step-size) The step-size is a constant $\alpha\in(0,1).$
Assumption 3.
(Unit bounded reward) We have
$\max_{(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}}\lvert r(s,a,s^{\prime})\rvert=R_{\text{max}}\leq 1.$
Assumption 4.
(Bounded initialization) The initial iterate $Q_{0}$ satisfies $\lVert Q_{0}\rVert_{\infty}\leq 1.$
Assumption 1 guarantees sufficient coverage of the state-action space and Assumption 3 and 4 are introduced without loss of generality and for simplicity of the analysis. For notational convenience, we define the following quantities that will be used throughout the paper.
@@
$d_{\mathrm{max}}\coloneq\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}d(s,a)\in(0,1).$
@@
$d_{\mathrm{min}}\coloneq\min_{(s,a)\in\mathcal{S}\times\mathcal{A}}d(s,a)\in(0,1).$
@@
$\rho\coloneq 1-\alpha d_{\mathrm{min}}(1-\gamma).$ (5)
Under Assumption 2, the decay rate satisfies $\rho\in(0,1)$. Throughout the paper, we will use the following compact notations for dynamical system representations:
[…]
and
$D_{a}=\mathrm{diag}\big(d(1,a),\ldots,d(|\mathcal{S}|,a)\big),\quad D=\mathrm{blkdiag}(D_{1},\ldots,D_{|\mathcal{A}|}).$
where $P_{a}=P(\cdot|a,\cdot)\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$, $Q(\cdot,a)\in\mathbb{R}^{|\mathcal{S}|}$ for $a\in\mathcal{A}$, and $R_{a}(s)\coloneqq\mathbb{E}[r(s,a,s^{\prime})|s,a]$. Here, $\mathrm{diag}(\cdot)$ denotes a diagonal matrix formed from its vector arguments, and $\mathrm{blkdiag}(\cdot)$ denotes a block-diagonal matrix whose diagonal blocks are the given matrices. Note that $P\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}|}$, $R,Q\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, and $D\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}||\mathcal{A}|}$. With this notation, the $Q$-function can be represented as a single stacked vector $Q\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ that enumerates all $Q(s,a)$ values for every $(s,a)\in\mathcal{S}\times\mathcal{A}$. Each entry $Q(s,a)$ can be expressed as $Q(s,a)=(e_{a}\otimes e_{s})^{T}Q$, where $e_{s}\in\mathbb{R}^{|\mathcal{S}|}$ and $e_{a}\in\mathbb{R}^{|\mathcal{A}|}$ denote the standard basis vectors, whose $s$-th and $a$-th components are equal to one and all other components are zero, respectively, and $\otimes$ denotes the Kronecker product. Under Assumption 2, the matrix $D$ is a nonsingular diagonal matrix with strictly positive diagonal elements. For any stochastic policy $\pi:\mathcal{S}\to\Delta_{|\mathcal{A}|}$, where $\Delta_{|\mathcal{A}|}$ denotes the probability simplex over $\mathcal{A}$, we define the matrix
$\Pi^{\pi}\coloneqq\begin{bmatrix}\pi(1)^{T}\otimes e_{1}^{T}\\ \pi(2)^{T}\otimes e_{2}^{T}\\ \vdots\\ \pi(\lvert S\rvert)^{T}\otimes e_{\lvert\mathcal{S}\rvert}^{T}\\ \end{bmatrix}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\times\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$ (6)
It is well known that $P\Pi^{\pi}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}||\mathcal{A}|}$ represents the transition probability matrix of state–action pairs under policy $\pi$. In the case of a deterministic policy $\pi:\mathcal{S}\to\mathcal{A}$, the stochastic policy can be equivalently expressed using a one-hot encoding vector $\vec{\pi}(s):=e_{\pi(s)}\in\Delta_{|\mathcal{A}|}$. The resulting action-transition matrix takes the same form as (6), with $\pi$ replaced by $\vec{\pi}$. For any $Q\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, we denote by $\pi_{Q}(s):=\arg\max_{a\in\mathcal{A}}Q(s,a)$ the greedy policy with respect to $Q$, and use the shorthand notation $\Pi_{Q}:=\Pi^{\pi_{Q}}$.
We recall a standard result ensuring that the $Q$-learning sequence remains bounded [23], which plays an important role in our analysis.
@@
$\lVert Q_{k}\rVert_{\infty}\leq Q_{\mathrm{max}}=\frac{\mathrm{max}{\{R_{\mathrm{max}},\mathrm{max}_{(s,a)\in\mathcal{S}\times\mathcal{A}}\|Q_{0}(s,a)\|_{\infty}\}}}{1-\gamma}.$
From Assumptions 3 and 4, we can easily see that $Q_{\max}\leq\frac{1}{1-\gamma}$.
@@ 4.1 Algorithm
$Q_{k+1}^{A}(s_{k},a_{k})$ $=Q_{k}^{A}(s_{k},a_{k})+\alpha_{k}\{r_{k+1}+\gamma Q_{k}^{A}(s_{k+1},\text{argmax}_{a\in A}Q_{k}^{B}(s_{k+1},a))-Q_{k}^{A}(s_{k},a_{k})\},$
$Q_{k+1}^{B}(s_{k},a_{k})$ $=Q_{k}^{B}(s_{k},a_{k})+\alpha_{k}\{r_{k+1}+\gamma Q_{k}^{B}(s_{k+1},\text{argmax}_{a\in A}Q_{k}^{A}(s_{k+1},a))-Q_{k}^{B}(s_{k},a_{k})\},$ (7)
@@
where $Q_{k}^{A}$ and $Q_{k}^{B}$ denote two separate estimators of the optimal action–value function $Q^{*}$ at iteration $k$. The pair $(s_{k},a_{k})\in\mathcal{S}\times\mathcal{A}$ represents the state–action pair sampled at time $k$, $r_{k+1}$ is the immediate reward observed after taking action $a_{k}$ at state $s_{k}$, and $s_{k+1}$ is the subsequent state. The scalar $\alpha_{k}>0$ denotes the step size at iteration $k$, and $\gamma\in(0,1)$ is the discount factor. The first difference between the original double $Q$-learning and SDQ is the role of each $Q$-estimator in the update. In the original double $Q$-learning, an optimal action is selected from the same $Q$-estimator, and it employs the other $Q$-estimator for bootstrapping. On the other hand, in the proposed version, an optimal action is selected from the other $Q$-estimator, and it employs the same $Q$-estimator for bootstrapping. This modification enables the use of the switching system framework from [12]. It overcomes the difficulty caused by the switched order of $Q^{A}_{k}$ and $Q^{B}_{k}$ in the original double $Q$-learning while retaining the advantage of reducing overestimation bias.
@@
The other difference is in the Bernoulli variable. Unlike the standard double $Q$-learning, which uses a Bernoulli variable for the $Q$-estimator selection, the modified version updates the two $Q$-estimators synchronously, which can potentially speed up the convergence. However, we note that our analysis can also include the Bernoulli random selection as in the original form without major changes in the finite-time error analysis. Besides, a potential issue that arises by eliminating the random $Q$-estimator selection is that if initially $Q^{A}_{0}=Q^{B}_{0}$, then $Q^{A}_{k}=Q^{B}_{k}$ for all $k\geq 0$. In this case, (4.1) is reduced to the standard $Q$-learning because in this case, $Q^{A}_{k}=Q^{B}_{k}$ for all $k\geq 0$. To bypass the issue for implementation, one simple approach is to randomly initialize $Q^{A}_{0}$ and $Q^{B}_{0}$ so that $Q^{A}_{0}\neq Q^{B}_{0}$.
@@
To demonstrate its effect, let us consider the example in Figure 1(left) (adopted from [1], Ch. 6.7). We consider an epsilon-greedy exploration with $\epsilon=0.1$, constant step‐size $\alpha=0.1$, and discount factor $\gamma=0.9$. The experiment consists of 1,000 independent runs, each comprising 500 training episodes. The initial QQ‐values for SDQ are uniformly sampled from [0.3,0.3][-0.3,0.3]. We also include a perturbed QQ‐learning variant, in which the QQ‐table is initialized by sampling from the same uniform distribution [0.3,0.3][-0.3,0.3]. As observed in the result, SDQ initially suffers from overestimation of QQ‐values due to the random initialization of its two estimators. However, this bias is quickly mitigated, and SDQ converges at a rate similar to that of original double QQ‐learning, as shown in Figure 1(right). Furthermore, both standard QQ‐learning and the perturbed QQ‐learning variant continue to exhibit overestimation, highlighting the efficacy of SDQ in mitigating this bias.
Figure 1: Left: An example from [1]. The episode always starts from the AA node. Taking the right action from the AA node results in zero reward, and the episode is terminated. Otherwise, taking the left action leads to state $B$, where the agent chooses one of 10 available actions. Executing any of these actions results in a reward sampled from a normal distribution with mean 0.1-0.1 and standard deviation 11. Then, the episode is terminated as well. Although $Q^{*}(A,\text{right})$ is zero and $Q^{*}(A,\text{left})$ is $0.1\gamma$, $Q$-learning favors left action because of maximization bias. Right: Comparison of experiment results: SDQ vs. double $Q$-learning vs. $Q$-learning vs. $Q$-learning (perturbed, with randomly initialized $Q$-values).
@@ 4.2 Experiment
We organize our evaluation into two complementary studies. First, we test the ability of SDQ to correct maximization bias in a simple stochastic 8×8 grid world where each step yields a stochastic reward. This environment makes overestimation bias clear and allows us to compare SDQ against both standard and perturbed $Q$-learning and double $Q$-learning. Here, the perturbed $Q$-learning variant refers to standard $Q$-learning with randomly initialized $Q$-values. Second, we demonstrate that SDQ converges faster than double $Q$-learning across three deterministic OpenAI Gym tasks, FrozenLake-v0, CliffWalking-v0, and Taxi-v3. All agents use the same epsilon-greedy exploration strategy, learning rate, and discount factor. Each agent is trained until its learning curve has fully stabilized. Together, these experiments show that SDQ not only nearly eliminates overestimation bias but also delivers consistent gains in convergence speed.
4.2.1 Grid World
We begin by evaluating SDQ in a simple stochastic grid‐world from Figure 2(left) designed to expose maximization bias. The agent occupies an 8×8 grid, starting in the lower‐left cell and seeking the upper‐right goal. Each non‐terminal transition yields a reward of 10-10 or +2+2 with equal probability, while entering the goal state grants +20+20 and immediately terminates the episode. All five algorithms, $Q$-learning, perturbed $Q$-learning, double $Q$-learning, perturbed double $Q$-learning, and SDQ, are run for 10,000 steps using epsilon-greedy where $\varepsilon(s)=1/\sqrt{n(s)}$ and $n(s)$ is the number of times state $s$ has been visited. The learning rate $\alpha_{k}(s,a)$ is chosen as a linear decay, $\alpha_{k}(s,a)=1/n_{k}(s,a)$. In the case of double $Q$-learning, the count $n_{k}(s,a)$ is set to $n_{k}^{A}(s,a)$ when updating $Q_{k}^{A}$, and to $n_{k}^{B}(s,a)$ when updating $Q_{k}^{B}$. The variables $n_{k}^{A}$ and $n_{k}^{B}$ respectively record how many times each state–action pair has been updated in the two value functions. The discount factor is set to $\gamma=0.95$, and QQ‐values are initialized by sampling uniformly from [0.3,0.3][-0.3,0.3], consistent with the setup in Figure 1. Results are averaged over 10 independent runs. Figure 2(middle) shows the average cumulative reward per step. SDQ achieves a marginally higher final return than the other methods. Figure 2(right) plots the maximum action-value at the start state, $\max_{a}Q_{k}(s_{0},a)$, against the true optimal value $\max_{a}Q^{*}(s_{0},a)$ (dashed line). Both standard and perturbed $Q$-learning clearly overestimate, while the two double $Q$-learning variants underestimate. SDQ stays closest to the optimum throughout. This result shows that it effectively corrects the overestimation bias.
4.2.2 FrozenLake, CliffWalking, and Taxi Environments
Next, we evaluate SDQ on three deterministic Gym tasks, FrozenLake-v0, CliffWalking-v0, and Taxi-v3, and show that it consistently converges faster than double $Q$-learning. In these experiments, we compare SDQ against original double $Q$-learning and a perturbed variant. We employ epsilon-greedy exploration with $\epsilon=0.1$, a constant learning rate $\alpha=0.01$, and a discount factor $\gamma=0.99$ for all algorithms. $Q$-estimators for SDQ are initialized randomly with the uniform distribution $[0,0.01]$, and for the perturbed double $Q$-learning variant we apply the same uniform initialization $[0,0.01]$ to both estimators. To account for the sparse rewards in FrozenLake-v0, we evaluate the average episodic reward after applying a moving window of size 100 for each episode, smoothing the reward signal since the agent only receives a +1+1 reward upon successful completion. Agents are trained for 10,000 episodes in FrozenLake-v0 to accommodate its sparse rewards, for 500 episodes in CliffWalking-v0, and for 30,000 episodes in Taxi-v3. These episode counts ensure that each environment’s learning curve has stabilized. For all experiments, the results are averaged over $30$ independent runs.
Figure 3 shows that SDQ achieves a modest but consistent improvement in convergence speed over both standard double $Q$-learning and its perturbed variant, suggesting this gain stems from its structural design rather than initialization alone. While the final returns are comparable to those of double $Q$-learning, SDQ generally reaches its steady-state performance faster, which is consistent with the theoretical insight on its improved stability.
SG
Figure 2: Left: 8×8 Grid world example. Middle: Average cumulative reward per step for each algorithm. Right: Evolution over time of the start‐state’s maximum action‐value.
@@ 4.2.2 FrozenLake, CliffWalking, and Taxi Environments
Figure 3: Comparison of experiment results: SDQ vs. double $Q$-learning vs. double $Q$-learning (perturbed, with randomly initialized $Q$-values).
Remark 4.1 (Applicability to complex environments).
Recent studies have extended the double-estimator framework of double $Q$-learning to more complex and high-dimensional domains, including continuous-control and real-world dynamic settings (e.g., [27, 28, 29]). These works demonstrate that the double-estimator structure remains a useful foundation for achieving stable and adaptive learning in complex environments. Unlike these approaches, which typically employ stochastic or alternating estimator updates, our SDQ adopts a deterministic coupling mechanism where both estimators are updated concurrently. This structural difference enables a tractable finite-time convergence analysis within a control-theoretic framework, clarifying the theoretical role of estimator coupling beyond its empirical advantages.